OpenStudy (dragon123mama):

what is the geometric mean of the pair of numbers? will give medal. not sure how to do this. 96 and 6

OpenStudy (sshayer):

G.M. of two numbers a and b is\[\sqrt{a*b}\]

OpenStudy (mathmale):

"geometric mean" is a concept I never learned during my school years. In situations like this, I look up the unknown term on Google (and suggest that you do the same). Here are some of my search results: https://www.google.com/search?sourceid=chrome-psyapi2&ion=1&espv=2&ie=UTF-8&q=geometric%20mean&oq=geometric%20mean&aqs=chrome..69i57j0l5.2481j0j7 The formula given you by sshayer is correct, and will "work," but there's more to "geometric mean" than that.

OpenStudy (dragon123mama):

when i solves it using that formula none of these answeres came up on the calculator. 1.24 2.32 3.51 4.16

OpenStudy (dragon123mama):

nvm i typed it in wrong. thanks that formula was a big help

OpenStudy (sshayer):

\[\sqrt{96*6}=\sqrt{16*6*6}=\sqrt{4*4*6*6}=?\]

OpenStudy (dragon123mama):

i already got the answer

OpenStudy (sshayer):

ok